# Optimal control under unknown intensity with Bayesian learning

### v1

@@ 3.3 Time regularity
Under Assumption 3.1, the value function is continuous.
In view of Proposition 3.2, all we have to prove is that the value function is continuous with respect to time. As it is assumed that the prior distribution is compactly supported in $[0,\lambda_{\max}]$, we recall that
@@
We start with a lemma. Let $\gamma\in{\mathcal{A}}$. We denote by $\varphi^{\gamma}_{t,s}(y)$ the solution of the ODE
@@
There exists a constant $C_{T}>0$ such that for all $t\leq s\leq T$, it holds that
@@
This follows from the Grönwall’s lemma applied to $s\mapsto\varphi^{\gamma}_{t,s}(y)-y$, using that $b$ is Lipschitz. ∎
Assume without loss of generalities that $s>t$ and that $s-t\leq 1$. To simplify the notations, we write $Y_{s}=Y^{t,y,\gamma}_{s},n_{s}=n+N_{s}-N_{t}$, $z_{s}=z+\int_{t}^{s}g(Y^{t,y,\gamma}_{u}\mathrm{d}u)$. Using Proposition 3.9,
@@
Because $v$ is bounded by $\frac{\lambda_{\max}^{2}}{4}$, and because $\mathbb{E}_{\mathbb{Q}}\langle L^{y,\gamma}_{t,s},m_{\mu}(n,z)\rangle\mathbbm{1}_{\{N_{s}>N_{t}\}}\leq\mathbb{P}(N_{s}>N_{t})\leq C(s-t)$, where under $\mathbb{P}$, $(N_{t})$ is a Poisson process of rate $\lambda_{\max}\lVert g\rVert_{\infty}$, we deduce that $|A^{\gamma}_{3}|\leq C(s-t)$ for any control $\gamma\in{\mathcal{A}}$. For the second term, we have using the results of Section 3.1:
@@
Using that $\lvert\gamma_{u}\rvert\leq 1+\gamma^{2}_{u}$, we finally deduce that for any control $\gamma\in{\mathcal{A}}$,
@@
Define the following sub-class of admissible control:
@@
Clearly, we can restrict to this class of control as:
@@
Fix $\epsilon>0$. Let $\gamma\in{\mathcal{A}}_{0}(t,y,z,n)$ be an $\epsilon$-optimal for (3.1). We have:
@@
Using the dominated convergence theorem, the left-hand-side is smaller that $3\epsilon$ when $s$ is sufficiently closes to $t$. Altogether, this proves the time continuity of the value function. ∎
@@ 3.4 The dynamic programming principle
We now state the general dynamical programming principle which allows from stopping times.
@@
Grant Assumptions 3.1. Let $\tau$ be a $({\mathcal{F}}^{N}_{t})$ stopping time such that, a.s., $\tau\in[t,T]$. Then
$v(t,y,z,n)=\inf_{\gamma\in{\mathcal{A}}}$ $\mathbb{E}_{\mathbb{Q}}\langle L^{y,\gamma}_{t,\tau},m_{\mu}(n,z)\rangle\left[\int_{t}^{\tau}\frac{\gamma^{2}_{u}}{2}\mathrm{d}u+v(\tau,Y^{t,y,\gamma}_{\tau},Z^{t,y,z,\gamma}_{\tau},n+N_{\tau}-N_{t})\right].$
@@
As the value function is continuous (Proposition 3.11), the proof is similar to the proof of Proposition 3.9, with a minor difference: the partition we consider is now a partition of $[0,T]\times\mathbb{R}\times\mathbb{R}_{+}$, to take into account the stopping times. Using Lemma 3.10, the proof is completed as before. ∎
@@ 4 Viscosity solution properties and characterization
In this section, to simplify the notations, we write $x=(y,z)\in\mathbb{R}\times\mathbb{R}_{+}$. With this convention, the value function is $v(t,x,n)$, $t\in[0,T]$ and $n\in\mathbb{N}$. By the previous section, we expect that the value function solves formally:
@@
with terminal condition
@@
$H(x,p):=\frac{1}{2}p_{1}^{2}-B(x)\cdot p,\quad B(x):=\begin{pmatrix}b(y)\\ g(y)\end{pmatrix}\quad\psi_{n}(x):=(0,z,n+1)$
and
@@
We denote by $D:=[0,T]\times\mathbb{R}\times\mathbb{R}_{+}$ and write $BUC(D)$ for the space of all bounded and uniformly continuous functions on $D$. Following [9], we consider the following notion of viscosity solutions.
@@
We say that a function $u:D\times\mathbb{N}\rightarrow\mathbb{R}$ is a viscosity sub-solution of (4.1) if for all $n\in\mathbb{N}$, $(t,x)\mapsto u(t,x,n)\in BUC(D)$ and if for all $(t_{\circ},x_{\circ},n_{\circ})\in D\times\mathbb{N}$, for all $\varphi\in C(D\times\mathbb{N})$ such that $\varphi$ is $C^{1}$ in a open neighborhood of $(t_{\circ},x_{\circ},n_{\circ})$ with $\max(u-\varphi)=(u-\varphi)(t_{\circ},x_{\circ},n_{\circ})=0$, it holds that if $t_{\circ}<T$:
@@
We say a function $w:D\times\mathbb{N}\rightarrow\mathbb{R}$ is a viscosity super-solution of (4.1) if for all $n\in\mathbb{N}$, $(t,x)\mapsto w(t,x,n)\in BUC(D)$ and if for all $(t_{\circ},x_{\circ},n_{\circ})\in D\times\mathbb{N}$, for all $\varphi\in C(D\times\mathbb{N})$ such that $\varphi$ is $C^{1}$ in a open neighborhood of $(t_{\circ},x_{\circ},n_{\circ})$ with $\min(w-\varphi)=(w-\varphi)(t_{\circ},x_{\circ},n_{\circ})=0$, it holds that if $t_{\circ}<T$:
$-\partial_{t}\varphi(t_{\circ},x_{\circ},n_{\circ})+H(x_{\circ},\partial_{x}\varphi(t_{\circ},x_{\circ},n_{\circ}))+\theta_{n_{\circ}}(x_{\circ})\left[w(t_{\circ},x_{\circ},n_{\circ})-w(t_{\circ},\psi_{n_{\circ}}(x_{\circ}))\right]\geq 0,$
Finally, we say that $u:D\times\mathbb{R}\rightarrow\mathbb{N}$ is a viscosity solution of (4.1) if it is both a viscosity sub-solution and a viscosity super-solution of this equation.
The main result of this section is
@@
Under Assumptions 3.1, the value function $v$ is the unique viscosity solution of (4.1) which satisfies the terminal condition (4.2).
We show successively that $v$ is a sub-solution, a super-solution of (4.1) and that a comparison theorem holds. We first need this crucial lemma:
@@
The process is integrable (since $\Lambda\in L^{1}(\mathbb{P})$) and $\mathcal{F}^{N}-$ adapted. Moreover, since
@@
$=\mathbb{E}\left[\int_{s}^{t}\mathbb{E}\left(\Lambda\mid\mathcal{F}_{u}^{N}\right)g(Y_{u}^{t,y,\gamma})\mathrm{d}u\mid\mathcal{F}_{s}^{N}\right],$
we obtain the result. ∎
4.1 Sub-solution property
In this section, we prove that the value function is a viscosity sub-solution of (4.1). To proceed, we use an equivalent reformulation of the viscosity solutions. We have, following [9]:
@@ 4.1 Sub-solution property
A function $u\in BUC(D)$ for all $n\in\mathbb{N}$ is a viscosity sub-solution of (4.1) if and only if for all $(t_{\circ},x_{\circ},n_{\circ})\in D\times\mathbb{N}$, for all $\varphi\in C(D\times\mathbb{N})$ such that $\varphi$ is $C^{1}$ in an open neighborhood of $(t_{\circ},x_{\circ},n_{\circ})$ and $\max(u-\varphi)=(u-\varphi)(t_{\circ},x_{\circ},n_{\circ})=0$, it holds that if $t_{0}<T$:
A similar statement holds for viscosity super-solution.
The only difference with the definition is in the zero order terms, where we have replaced the sub-solution $u$ by the test function $\varphi$. For the proof, see [9, Lem. 2.1]. We now prove that:
Lemma 4.5.
The value function is a viscosity sub-solution of (4.1).
@@
With $x_{\circ}:=(y_{\circ},z_{\circ})$, let $(t_{\circ},x_{\circ},n_{\circ})\in D\times\mathbb{N}$ and let $\varphi:D\times\mathbb{N}\rightarrow\mathbb{R}$ a test function as in Lemma 4.4. Let $\gamma\in\mathbb{R}$ be a constant and deterministic control. Let $n_{s}:=n_{\circ}+N_{s}-N_{t_{\circ}}$ and $X_{s}^{t_{\circ},x_{\circ},\gamma}=(Y_{s}^{t_{\circ},y_{\circ},\gamma},Z_{s}^{t_{\circ},y_{\circ},z_{\circ},\gamma}),t_{\circ}\leq s\leq T$. Consider $T_{1}$ the time of the first jump:
$T_{1}=\inf\{s\geq t_{\circ},N_{s}\neq N_{s-}\}.$
As $\gamma$ is a valid control and $s\wedge T_{1}$ a stopping time we have by the dynamical programming principle 3.13:
@@
Because $\varphi\geq v$, we deduce that
@@
where $\mathrm{d}\tilde{N}_{u}=\mathrm{d}N_{u}-\theta_{n_{u}}(X_{u}^{t_{\circ},x_{\circ},\gamma})\mathrm{d}u=\mathrm{d}N_{u}-\mathbb{E}[\Lambda~{}|~{}{\mathcal{F}}^{N}_{u}]g(Y_{u}^{t_{\circ},y_{\circ},\gamma})\mathrm{d}u$. Taking the expectation with respect to $\mathbb{E}_{\mathbb{P}}$ and using Lemma 4.3, we deduce that
@@
Note that because $u<T_{1}$, it holds that $n_{u}=n_{\circ}$, and similarly, $X_{u}^{t_{\circ},x_{\circ},\gamma}$ is a deterministic function. We choose $s=t_{\circ}+\frac{1}{k}$, $k\in\mathbb{N}^{*}$ large enough such that $(u,X_{u}^{t_{\circ},x_{\circ},\gamma})$ belongs to the open neighborhood for which the test function $\varphi$ is $C^{1}$, for all $u\leq s\wedge T_{1}$. By applying first the dominated convergence theorem and finally applying the mean value theorem, we deduce that:
@@
As this is true for all $\gamma$, we choose the optimal value $\gamma=-\partial_{y}\varphi(t_{\circ},x_{\circ},n_{\circ})$ and obtain exactly the required inequality of Lemma 4.4. ∎
4.2 Super-solution property
@@
The value function is a viscosity super-solution of (4.1).
@@
We rely similarly on Lemma 4.4. In addition, we note that we can restrict to the class of test functions $\varphi$ for which the minimum is strict, namely $(v-\varphi)(t_{\circ},x_{\circ},n_{\circ})=0$ and:
@@
We consider $\varphi$ such a test function. We work towards a contradiction and assume that
@@
    There exists $r>0$ small enough such that
@@
    holds on $B(t_{\circ},x_{\circ},n_{\circ};r)$ which is the ball of center $(t_{\circ},x_{\circ},n_{\circ})$ and of radius $r$. We introduce
@@
    which is bounded. By (4.4), there exists $\eta>0$ such that
@@
    Let $\gamma$ be an arbitrary control and $\theta_{\gamma}$ be the first exit time of $B(t_{\circ},x_{\circ},n_{\circ};r)$. By Itô’s formula, we obtain:
@@
    where $\widetilde{N}$ is the martingale process introduced in Lemma 4.3. Note that (4.5) can be rewritten as
@@ 4.3 Comparison theorem
Finally, we provide a comparison principle for the viscosity solutions of (4.1). The main result of this section is:
@@
Let $u$ be a viscosity sub-solution and let $w$ be a viscosity super-solution of the PDE (4.1), in the sense of Definition 4.1. Assume that $u(T,x,n)\leq w(T,x,n)$ for all $(x,n)\in(\mathbb{R}\times\mathbb{R}_{+})\times\mathbb{N}$. Then
@@
In particular, there is at most a bounded and uniformly continuous viscosity solution of (4.1) which satisfies the boundary condition $v(T,x,n)=\kappa\Psi_{n}(x)$.
@@
We proceed towards a contradiction and assume that $\theta:=\sup_{D\times\mathbb{N}}u-w>0$.
@@
    Let $\epsilon,\delta,\beta,\alpha>0$. We consider the auxiliary function
@@
    where $\lVert x\rVert:=\sqrt{y^{2}+z^{2}}$ with $x=(y,z)$ and:
@@
    For all $\epsilon,\delta,\beta,\alpha$ small enough, it holds that
@@
    In addition, as $u,w$ are bounded and $\delta>0,\alpha>0$, the $\sup$ is reached for some point $\bar{t},\bar{x},\bar{t}^{\prime},\bar{x}^{\prime},\bar{n}$.
@@
    From $\Phi_{\bar{n}}(\bar{t},\bar{x},\bar{t}^{\prime},\bar{x}^{\prime})\geq 0$ and $u,w$ bounded, we deduce that for $C=\lVert u\rVert_{\infty}+\lVert w\rVert_{\infty}$:
@@
    In addition, from $\Phi_{\bar{n}}(\bar{t},\bar{x},\bar{t^{\prime}},\bar{x^{\prime}})\geq\Phi_{\bar{n}}(\bar{t^{\prime}},\bar{x^{\prime}},\bar{t^{\prime}},\bar{x^{\prime}})$, we obtain:
@@
    For $\rho\geq 0$, let $D_{\rho}=\{(t,x),(t^{\prime},x^{\prime})\in D^{2}:|t-t^{\prime}|^{2}+\lVert x-x^{\prime}\rVert_{2}^{2}\leq\rho\}$, and
@@
    We deduce from $\lVert\bar{x}-\bar{x^{\prime}}\rVert^{2}+|\bar{t}-\bar{t^{\prime}}|^{2}\leq 2C\epsilon$ and from $\lvert\lVert\bar{x}\rVert-\lVert\bar{x^{\prime}}\rVert\rvert\leq\lVert\bar{x}-\bar{x^{\prime}}\rVert$ that:
@@
    As $(t,x)\mapsto u(t,x,n)$ is uniformly continuous and bounded, the function $m^{\alpha}_{u}:\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}$ is bounded, continuous and $m^{\alpha}_{u}(0+)=m^{\alpha}_{u}(0)=0$. A similar statement holds for $m^{\alpha}_{w}$.
@@
    Similarly, if $\bar{t^{\prime}}=T$, we have:
@@
    Assume that both $\bar{t}<T$ and $\bar{t^{\prime}}<T$. Consider $\varphi_{u}(t,x):=\frac{1}{2\epsilon}\lVert x-\bar{x^{\prime}}\rVert^{2}+\frac{1}{2\epsilon}|t-\bar{t^{\prime}}|^{2}+\delta F(\lVert x\rVert)$. It holds that $(\bar{t},\bar{x})\in\text{argmax}(u-\varphi_{u})$. So, by the property of viscosity subsolution of $u$, we have:
@@
    Similarly, consider $\varphi_{v}(t^{\prime},x^{\prime})=-\frac{1}{2\epsilon}\lVert\bar{x}-x^{\prime}\rVert^{2}-\frac{1}{2\epsilon}|\bar{t}-t^{\prime}|^{2}-\delta F(\lVert x^{\prime}\rVert)+\beta(t^{\prime}-T)$. As $(\bar{t^{\prime}},\bar{x^{\prime}})\in\text{argmin}(w-\varphi_{v})$, we have
@@
    Altogether, combining the two inequalities, we obtain $\beta\leq\Delta_{1}+\Delta_{2}+\Delta_{3}$ where:
@@
    We start with $\Delta_{1}$. We use the explicit shape of $H(x,p)=\frac{1}{2}p_{1}^{2}-B(x)\cdot p$. We have
@@
    we retain that $\Delta_{1}\leq k(\alpha,\epsilon,\delta)$. For $\Delta_{2}$, from Lemma 2.15 and since $g\in C^{1}$, we have:
    $L:=\sup_{n\in\mathbb{N}}\lVert\nabla\theta_{n}\rVert_{\infty}<+\infty.$
    We deduce that:
@@
    Finally, we treat $\Delta_{3}$. Let $\bar{x}^{*}=(0,\bar{x}_{2})$ and $\bar{x}^{\prime*}=(0,\bar{x}^{\prime}_{2})$. Using that

### v2

@@ 3.3 Time regularity
Under Assumption 3.1, the value function $v$ is continuous.
Given Proposition 3.2, it suffices to prove that the value function is continuous with respect to time. Since the prior distribution is compactly supported in $[0,\lambda_{\max}]$, we recall that:
@@
We begin with a preliminary lemma. For $\gamma\in{\mathcal{A}}$, let $\varphi^{\gamma}_{t,s}(y)$ denote the solution to the ordinary differential equation:
@@
There exists a constant $C_{T}>0$ such that, for all $t\leq s\leq T$:
@@
The result follows by applying Grönwall’s lemma to $s\mapsto\varphi^{\gamma}_{t,s}(y)-y$, using the Lipschitz property of $b$. ∎
Assume without loss of generality that $s>t$ and $s-t\leq 1$. To simplify notation, define:
$Y_{s}:=Y^{t,y,\gamma}_{s},\quad n_{s}:=n+N_{s}-N_{t},\quad z_{s}:=z+\int_{t}^{s}g(Y^{t,y,\gamma}_{u}\mathrm{d}u).$
Using Proposition 3.9, we write:
@@
Since $v$ is bounded by $\frac{\lambda_{\max}^{2}}{4}$, and $\mathbb{E}_{\mathbb{Q}}\langle L^{y,\gamma}_{t,s},m_{\mu}(n,z)\rangle\mathbbm{1}_{\{N_{s}>N_{t}\}}\leq\mathbb{P}(\tilde{N}_{s}>\tilde{N}_{t})\leq C(s-t)$, where under $\mathbb{P}$, $(\tilde{N}_{t})$ is a Poisson process with intensity $\lambda_{\max}\lVert g\rVert_{\infty}$, it follows that $|A^{\gamma}_{3}|\leq C(s-t)$ for any control $\gamma\in{\mathcal{A}}$. For the second term, using the regularity results from Section 3.1:
@@
Since $\lvert\gamma_{u}\rvert\leq 1+\gamma^{2}_{u}$, we deduce that, for any $\gamma\in{\mathcal{A}}$:
@@
Define the subclass of admissible controls:
@@
Clearly, we can restrict to this class since:
@@
Fix $\epsilon>0$. Let $\gamma\in{\mathcal{A}}_{0}(t,y,z,n)$ be an $\epsilon$-optimal for (3.1). Then:
@@
By the dominated convergence theorem, the right-hand side is less than $3\epsilon$ when $s$ is sufficiently close to $t$. Thus, the value function is continuous in time. ∎
@@ 3.4 The dynamic programming principle
We now establish the general dynamic programming principle, which extends to stopping times.
@@
Under Assumption 3.1, let $\tau$ be an $({\mathcal{F}}^{N}_{t})$ stopping time such that, a.s. $\tau\in[t,T]$. Then:
$v(t,y,z,n)=\inf_{\gamma\in{\mathcal{A}}_{t}}$ $\mathbb{E}_{\mathbb{Q}}\langle L^{y,\gamma}_{t,\tau},m_{\mu}(n,z)\rangle\left[\int_{t}^{\tau}\frac{\gamma^{2}_{u}}{2}\mathrm{d}u+v(\tau,Y^{t,y,\gamma}_{\tau},Z^{t,y,z,\gamma}_{\tau},n+N_{\tau}-N_{t})\right].$
@@
Given the continuity of the value function (Proposition 3.11), the proof follows similarly to that of Proposition 3.9, with a slight modification: the partition is now defined on $[0,T]\times\mathbb{R}\times\mathbb{R}_{+}$, to account for stopping times. Using Lemma 3.10, the proof is completed as in the deterministic case. ∎
@@ 4 Viscosity solution properties and characterization
In this section, to simplify notation, we let $x=(y,z)\in\mathbb{R}\times\mathbb{R}_{+}$. With this convention, the value function is denoted by $v(t,x,n)$, where $t\in[0,T]$ and $n\in\mathbb{N}$. Based on the results of the previous section, we expect that the value function formally satisfies:
@@
with the terminal condition:
@@
$H(x,p):=\frac{1}{2}p_{1}^{2}-B(x)\cdot p,\quad B(x):=\begin{pmatrix}b(y)\\ g(y)\end{pmatrix},\quad\psi_{n}(x):=(0,z,n+1)$
and:
@@
We denote $D:=[0,T]\times\mathbb{R}\times\mathbb{R}_{+}$ and let $BUC(D)$ denote the space of bounded and uniformly continuous functions on $D$. Following [13], we consider the following notion of viscosity solutions.
@@
A function $u:D\times\mathbb{N}\rightarrow\mathbb{R}$ is a viscosity subsolution of (4.1) if, for all $n\in\mathbb{N}$, the mapping $(t,x)\mapsto u(t,x,n)\in BUC(D)$, and for all $(t_{\circ},x_{\circ},n_{\circ})\in D\times\mathbb{N}$, for all functions $\varphi\in C(D\times\mathbb{N})$ that are $C^{1}$ in an open neighborhood of $(t_{\circ},x_{\circ},n_{\circ})$ and satisfy $\max(u-\varphi)=(u-\varphi)(t_{\circ},x_{\circ},n_{\circ})=0$, it holds that if $t_{\circ}<T$:
@@
A function $w:D\times\mathbb{N}\rightarrow\mathbb{R}$ is a viscosity supersolution of (4.1) if, for all $n\in\mathbb{N}$, the mapping $(t,x)\mapsto w(t,x,n)\in BUC(D)$, and for all $(t_{\circ},x_{\circ},n_{\circ})\in D\times\mathbb{N}$, for all functions $\varphi\in C(D\times\mathbb{N})$ that are $C^{1}$ in an open neighborhood of $(t_{\circ},x_{\circ},n_{\circ})$ and satisfy $\min(w-\varphi)=(w-\varphi)(t_{\circ},x_{\circ},n_{\circ})=0$, it holds that if $t_{\circ}<T$:
$-\partial_{t}\varphi(t_{\circ},x_{\circ},n_{\circ})+H(x_{\circ},\partial_{x}\varphi(t_{\circ},x_{\circ},n_{\circ}))+\theta_{n_{\circ}}(x_{\circ})\left[w(t_{\circ},x_{\circ},n_{\circ})-w(t_{\circ},\psi_{n_{\circ}}(x_{\circ}))\right]\geq 0.$
Finally, a function $u:D\times\mathbb{N}\rightarrow\mathbb{R}$ is a viscosity solution of (4.1)if it is both a viscosity subsolution and a viscosity supersolution of this equation.
We shall use an equivalent reformulation of viscosity solutions. Following [13], we have:
Lemma 4.2.
A function $u:D\times\mathbb{N}\mapsto\mathbb{R}$ with $(t,x)\mapsto u(t,x,n)\in BUC(D)$ for all $n\in\mathbb{N}$ is a viscosity subsolution of (4.1) if and only if, for all $(t_{\circ},x_{\circ},n_{\circ})\in D\times\mathbb{N}$, and for all functions $\varphi\in C(D\times\mathbb{N})$ that are $C^{1}$ in an open neighborhood of $(t_{\circ},x_{\circ},n_{\circ})$ and satisfy $\max(u-\varphi)=(u-\varphi)(t_{\circ},x_{\circ},n_{\circ})=0$, it holds that if $t_{0}<T$:
A similar statement holds for viscosity supersolutions.
The only difference from Definition 4.1 is in the zero-order term, where the subsolution $u$ is replaced by the test function $\varphi$. For the proof, see [13, Lem. 2.1].
The main result of this section is the following.
@@
Under Assumption 3.1, the value function $v$ is the unique viscosity solution of (4.1) that satisfies the terminal condition (4.2).
We prove successively that $v$ is a viscosity subsolution and a viscosity supersolution of (4.1), and that a comparison theorem holds. We first establish the following crucial lemma.
@@
The process is integrable (since $\Lambda\in L^{1}(\mathbb{P})$) and $\mathcal{F}^{N}-$ adapted. Moreover:
@@
$=\mathbb{E}\left[\int_{s}^{t}\mathbb{E}\left(\Lambda\mid\mathcal{F}_{u}^{N}\right)g(Y_{u}^{t,y,\gamma})\mathrm{d}u\mid\mathcal{F}_{s}^{N}\right].$
Thus, we obtain the result. ∎
4.1 Subsolution property
In this section we prove:
@@ 4.1 Subsolution property
The value function $v$ is a viscosity subsolution of (4.1).
@@
Let $x_{\circ}:=(y_{\circ},z_{\circ})$, $(t_{\circ},x_{\circ},n_{\circ})\in D\times\mathbb{N}$, and let $\varphi:D\times\mathbb{N}\rightarrow\mathbb{R}$ be a test function as in Lemma 4.2. Consider a constant deterministic control $\gamma\in\mathbb{R}$. Define $n_{s}:=n_{\circ}+N_{s}-N_{t_{\circ}}$ and $X_{s}^{t_{\circ},x_{\circ},\gamma}=(Y_{s}^{t_{\circ},y_{\circ},\gamma},Z_{s}^{t_{\circ},y_{\circ},z_{\circ},\gamma})$, for $t_{\circ}\leq s\leq T$. Let $T_{1}:=\inf\{s\geq t_{\circ},N_{s}\neq N_{s-}\}$ denote the time of the first jump. Since $\gamma$ is a valid control and $s\wedge T_{1}$ is a stopping time, the dynamic programming principle 3.13 implies:
@@
Since $\varphi\geq v$, it follows that:
@@
where $\mathrm{d}\tilde{N}_{u}=\mathrm{d}N_{u}-\theta_{n_{u}}(X_{u}^{t_{\circ},x_{\circ},\gamma})\mathrm{d}u=\mathrm{d}N_{u}-\mathbb{E}[\Lambda~{}|~{}{\mathcal{F}}^{N}_{u}]g(Y_{u}^{t_{\circ},y_{\circ},\gamma})\mathrm{d}u$. Taking the expectation with respect to $\mathbb{E}_{\mathbb{P}}$ and using Lemma 4.4, we obtain:
@@
Since $u<T_{1}$, we have $n_{u}=n_{\circ}$, and similarly, $X_{u}^{t_{\circ},x_{\circ},\gamma}$ is a deterministic function. Choose $s=t_{\circ}+\frac{1}{k}$, $k\in\mathbb{N}^{*}$, sufficiently large so that $(u,X_{u}^{t_{\circ},x_{\circ},\gamma})$ remains in the open neighborhood where $\varphi$ is $C^{1}$, for all $u\leq s\wedge T_{1}$. Applying the dominated convergence theorem and then the mean value theorem, we deduce:
@@
Optimizing over $\gamma$ by choosing $\gamma=-\partial_{y}\varphi(t_{\circ},x_{\circ},n_{\circ})$, we obtain the inequality required by Lemma 4.2. ∎
4.2 Supersolution property
@@
The value function $v$ is a viscosity supersolution of (4.1).
@@
We rely on Lemma 4.2. Additionally, we may restrict to test functions $\varphi$ for which the minimum is strict, i.e., $(v-\varphi)(t_{\circ},x_{\circ},n_{\circ})=0$ and:
@@
Consider $\varphi$ such a test function $\varphi$. Suppose, for contradiction, that:
@@
    There exists $r>0$ sufficiently small such that:
@@
    holds on $B(t_{\circ},x_{\circ},n_{\circ};r)$ the ball centered at $(t_{\circ},x_{\circ},n_{\circ})$ with radius $r$. Define:
@@
    which is bounded. By (4.4), there exists $\eta>0$ such that:
@@
    Let $\gamma$ be an arbitrary control, and let $\theta_{\gamma}$ be the first exit time of $B(t_{\circ},x_{\circ},n_{\circ};r)$. By Itô’s formula:
@@
    where $\widetilde{N}$ is the martingale process introduced in Lemma 4.4. From (4.5) we have:
@@ 4.3 Comparison theorem
We now establish a comparison principle for viscosity solutions of (4.1). The main result of this section is:
@@
Let $u$ be a viscosity subsolution and let $w$ be a viscosity supersolution of the PDE (4.1) in the sense of Definition 4.1. Assume that $u(T,x,n)\leq w(T,x,n)$ for all $(x,n)\in(\mathbb{R}\times\mathbb{R}_{+})\times\mathbb{N}$. Then:
@@
In particular, there exists at most one bounded and uniformly continuous viscosity solution of (4.1) that satisfies the terminal condition $v(T,x,n)=\kappa\Psi_{n}(x)$.
@@
Suppose, for contradiction, $\theta:=\sup_{D\times\mathbb{N}}u-w>0$.
@@
    Let $\epsilon,\delta,\beta,\alpha>0$. Consider the auxiliary function:
@@
    where $\lVert x\rVert:=\sqrt{y^{2}+z^{2}}$ for $x=(y,z)$ and:
@@
    For $\epsilon,\delta,\beta,\alpha$ sufficiently small, it holds that:
@@
    Since $u$ and $w$ are bounded and $\delta>0,\alpha>0$, the supremum is attained at some point $\bar{t},\bar{x},\bar{t}^{\prime},\bar{x}^{\prime},\bar{n}$.
@@
    From $\Phi_{\bar{n}}(\bar{t},\bar{x},\bar{t}^{\prime},\bar{x}^{\prime})\geq 0$ and the boundedness of $u$ and $w$, with $C=\lVert u\rVert_{\infty}+\lVert w\rVert_{\infty}$, we have:
@@
    From $\Phi_{\bar{n}}(\bar{t},\bar{x},\bar{t^{\prime}},\bar{x^{\prime}})\geq\Phi_{\bar{n}}(\bar{t^{\prime}},\bar{x^{\prime}},\bar{t^{\prime}},\bar{x^{\prime}})$, it follows:
@@
    For $\rho\geq 0$, define: $D_{\rho}=\{(t,x),(t^{\prime},x^{\prime})\in D^{2}:|t-t^{\prime}|^{2}+\lVert x-x^{\prime}\rVert_{2}^{2}\leq\rho\}$, and:
@@
    Since $\lVert\bar{x}-\bar{x^{\prime}}\rVert^{2}+|\bar{t}-\bar{t^{\prime}}|^{2}\leq 2C\epsilon$ and $\lvert\lVert\bar{x}\rVert-\lVert\bar{x^{\prime}}\rVert\rvert\leq\lVert\bar{x}-\bar{x^{\prime}}\rVert$, we deduce:
@@
    As $(t,x)\mapsto u(t,x,n)$ is uniformly continuous and bounded, $m^{\alpha}_{u}:\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}$ is bounded, continuous, and $m^{\alpha}_{u}(0+)=m^{\alpha}_{u}(0)=0$. A similar statement holds for $m^{\alpha}_{w}$.
@@
    Similarly, if $\bar{t^{\prime}}=T$, then:
@@
    If $\bar{t}<T$ and $\bar{t^{\prime}}<T$, consider:
    $\varphi_{u}(t,x):=\frac{1}{2\epsilon}\lVert x-\bar{x^{\prime}}\rVert^{2}+\frac{1}{2\epsilon}|t-\bar{t^{\prime}}|^{2}+\delta F(\lVert x\rVert).$
    Since $(\bar{t},\bar{x})\in\text{argmax}(u-\varphi_{u})$, the viscosity subsolution property of $u$ implies:
@@
    Similarly, consider:
    $\varphi_{v}(t^{\prime},x^{\prime})=-\frac{1}{2\epsilon}\lVert\bar{x}-x^{\prime}\rVert^{2}-\frac{1}{2\epsilon}|\bar{t}-t^{\prime}|^{2}-\delta F(\lVert x^{\prime}\rVert)+\beta(t^{\prime}-T).$
    Since $(\bar{t^{\prime}},\bar{x^{\prime}})\in\text{argmin}(w-\varphi_{v})$, the viscosity supersolution property of implies:
@@
    Combining these inequalities, we obtain $\beta\leq\Delta_{1}+\Delta_{2}+\Delta_{3}$ where:
@@
    For $\Delta_{1}$ using $H(x,p)=\frac{1}{2}p_{1}^{2}-B(x)\cdot p$, we have:
@@
    we retain that $\Delta_{1}\leq k(\alpha,\epsilon,\delta)$. For $\Delta_{2}$, since $g\in C^{1}$ and by Lemma 2.15, $L:=\sup_{n\in\mathbb{N}}\lVert\nabla\theta_{n}\rVert_{\infty}<+\infty$, so:
@@
    For $\Delta_{3}$, let $\bar{x}^{*}=(0,\bar{x}_{2})$ and $\bar{x}^{\prime*}=(0,\bar{x}^{\prime}_{2})$. Since: